OpenStudy (thatbishtay17):

A scientist is measuring the amount of radioactive material in an unknown substance. When he begins measuring, there are 6.2 grams of radioactive substance. Twelve days later, there are 6.16 grams. After 19 days, there are 6.13 grams. After 46 days, there are 5.92 grams. Assuming that the decay is exponential, find the equation that determines the number of grams remaining after x days. Help!!! A. y = 6(0.92)x B. y = 6.15(0.98)x C. y = 6.23(0.999)x D. y = 6(0.999)x

OpenStudy (thatbishtay17):

@karatechopper

OpenStudy (karatechopper):

Any idea of where to start?

OpenStudy (thatbishtay17):

No clue honestly

OpenStudy (karatechopper):

are all these numbers being multiplied together or is there an exponent anywhere

OpenStudy (thatbishtay17):

Not that I can see.

OpenStudy (karatechopper):

This seems to be an exponential decay problem in which I woudl normally see "e" so I'm a bit lost here

OpenStudy (wolf1728):

I have found a formula - this should help: The second formula down is what you should use. I'll help you too.

OpenStudy (karatechopper):

ah so there is log and exponents

OpenStudy (wolf1728):

half life = elapsed time * log(2) / [log (bgng amt/ending amt)] start 6.2 g 12 days 6.16 g half life = 12 * .3010 / log (6.2 / 6.16) half life =3.612 / 0.0028109773366 half life = 1,284.96 days Let's see if that is correct There's a half life calculator here: http://www.1728.org/halflife.htm

OpenStudy (wolf1728):

Calculator says 1,285.1 For 19 days and 6.13 grams half life = 1,160 days For 46 days and 5.92 grams half life = 690 days something must be wrong with the data. Is that typed correctly?

OpenStudy (wolf1728):

Hey welcome back!!

OpenStudy (wolf1728):

As far as the formula, I make it: y = 6.2 * .99945^days

OpenStudy (wolf1728):

Is that problem stated without any typos?